Circular No. 8171

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gs IN NGC 936 R. Evans, Hazelbrook, N.S.W., reports his visual discovery of a su-

pernova (mag 14.0) on July 29.75 UT with a 0.31-m reflector, noting it to be located ∼ 20′′ southeast of the nucleus of NGC 936 (α = 2^{h}27^{m}.6, δ = −1^{o}09′, equinox 2000.0); nothing was detected to mag 15 by Evans in a search of this galaxy on July 3.75. Evans again found SN 2003gs at m_{v} = 13.8 on July 30.75. H. Yamaoka, Kyushu University, reports red mag ∼ 13.2 and the following precise position for SN 2003gs from an unfiltered CCD image taken on July 30.767 by T. Dobosz, Sydney, N.S.W.: α = 2^{h}27^{m}38^{s}.36, δ = −1^{o}09′35′′.4 (equinox 2000.0; fourteen GSC 2.2 reference stars; fitting error ∼ 0′′.3 in each coordinate), which is 13′′.4 east and 14′′.6 south of the center of NGC 936. N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports that a spectrogram (range 350–900 nm; resolution 2 nm) of SN 2003gs was obtained by P. Candia and M. Stritzinger on July 30.4 with the facility spectrograph on the CTIO 1.5-m telescope. The spectrum is peculiar and is roughly consistent with a type-Ia supernova perhaps 1 week after maximum light; the Si II lines are redshifted by ∼ 11000 km/s and are weak for this phase of the supernova. K. Krisciunas, CTIO, reports the following preliminary magnitudes for SN 2003gs from data obtained by J. Espinoza and D. Depoy with the CTIO 1.3-m telescope on July 30.40: U ∼ 14.70, B ∼ 14.05, V ∼ 13.40.

NO SUPERNOVAE 1998ev, 1999cu, AND 1999dc A. Gal-Yam and D. Maoz, Tel Aviv University, write that reanalysis

(using improved image-subtraction software) of numerous unfiltered CCD images of the field of the galaxy cluster Abell 136, obtained with the Wise Observatory 1-m telescope, shows that the the source reported as SN 1998ev (cf. IAUC 7055) is probably not variable. Similar reanalysis of images of the field of the galaxy cluster Abell 1697 shows that the source reported as SN 1999cu (cf. IAUC 7210) is indeed variable, although certainly not a supernova, with a temporal behavior consistent with that of an active galactic nucleus (AGN). R. Chornock and A. V. Filippenko, University of California, Berkeley, report that inspection of a CCD spectrum (range 320– 930 nm), obtained on 2003 June 29 UT with the Keck-I telescope, shows that the host galaxy of the object 1999dc (cf. IAUC 7227) is, in fact, an AGN at a redshift of 0.791; Gal-Yam and Maoz therefore conclude that 1999dc was probably not a supernova, either.

2003 July 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green